# Random Policy Enables In-Context Reinforcement Learning within Trust Horizons

### v2

@@ 5 Conclusion
Appendix A Related Work
A.1 Offline Reinforcement Learning
A.2 Autoregressive-Supervised Decision Making
A.3 In-Context Reinforcement Learning
Appendix B Omitted Proofs
B.1 Proof of Theorem 1
Theorem 1 (MAB). Let Assumption 1 hold. Denote by $s_{q}$ and $a^{*}$ the singleton state and optimal arm in the MAB problem. Consider a random policy $\pi$. Suppose that each arm has been selected $N$ times ($N\in{\mathbb{N}}_{+}$, trust horizon) under $\pi$. With probability at least $1-\delta$, it holds that
$\frac{1}{N}\sum_{i=1}^{N}r(s_{q},a^{*}\mid\pi)\geq\max_{a\in A\setminus\{a^{*}\}}\frac{1}{N}\sum_{i=1}^{N}r(s_{q},a\mid\pi),$ (17)
when the trust horizon $N$ satisfies
@@ B.1 Proof of Theorem 1
$\frac{1}{N}\sum_{i=1}^{N}r(s_{q},a^{*}\mid\pi)\geq{\mathbb{E}}[r(s_{q},a^{*})]-\epsilon_{1},$ (21)
$\frac{1}{N}\sum_{i=1}^{N}r(s_{q},a\mid\pi)\leq{\mathbb{E}}[r(s_{q},a)]+\epsilon_{2}.$ (22)
We note that the two inequalities above are the sufficient but not necessary conditions for $\frac{1}{N}\sum_{i=1}^{N}r(s_{q},a^{*}\mid\pi)\geq\frac{1}{N}\sum_{i=1}^{N}r(s_{q},a\mid\pi)$ to hold.
@@
$P\left(\frac{1}{N}\sum_{i=1}^{N}r(s_{q},a^{*}\mid\pi)\geq\frac{1}{N}\sum_{i=1}^{N}r(s_{q},a\mid\pi)\right)$
$\geq P\left(\frac{1}{N}\sum_{i=1}^{N}r(s_{q},a^{*}\mid\pi)\geq{\mathbb{E}}[r(s_{q},a^{*})]-\epsilon_{1},\frac{1}{N}\sum_{i=1}^{N}r(s_{q},a\mid\pi)\leq{\mathbb{E}}[r(s_{q},a)]+\epsilon_{2}\right)$ (23)
$=P\left(\frac{1}{N}\sum_{i=1}^{N}r(s_{q},a^{*}\mid\pi)\geq{\mathbb{E}}[r(s_{q},a^{*})]-\epsilon_{1}\right)\cdot P\left(\frac{1}{N}\sum_{i=1}^{N}r(s_{q},a\mid\pi)\leq{\mathbb{E}}[r(s_{q},a)]+\epsilon_{2}\right),$ (24)
@@
$P\left(\frac{1}{N}\sum_{i=1}^{N}r(s_{q},a\mid\pi)-{\mathbb{E}}[r(s_{q},a)]\leq\epsilon_{2}\right)\geq 1-\exp\left(-\frac{2N\epsilon_{2}^{2}}{(B-(-B))^{2}}\right),$ (25)
@@
$P\left(\frac{1}{N}\sum_{i=1}^{N}r(s_{q},a\mid\pi)-{\mathbb{E}}[r(s_{q},a)]\leq\epsilon_{2}\right)\geq 1-\exp\left(-\frac{N\epsilon_{2}^{2}}{2B^{2}}\right).$ (26)
@@
$P\left(\frac{1}{N}\sum_{i=1}^{N}r(s_{q},a^{*}\mid\pi)-{\mathbb{E}}[r(s_{q},a^{*})]\geq-\epsilon_{1}\right)\geq 1-\exp\left(-\frac{2N\epsilon_{1}^{2}}{(B-(-B))^{2}}\right),$ (27)
@@
$P\left(\frac{1}{N}\sum_{i=1}^{N}r(s_{q},a^{*}\mid\pi)-{\mathbb{E}}[r(s_{q},a^{*})]\geq-\epsilon_{1}\right)\geq 1-\exp\left(-\frac{N\epsilon_{1}^{2}}{2B^{2}}\right).$ (28)
@@
$P\left(\frac{1}{N}\sum_{i=1}^{N}r(s_{q},a^{*}\mid\pi)\geq\frac{1}{N}\sum_{i=1}^{N}r(s_{q},a\mid\pi)\right)$
$\geq\left(1-\exp\left(-\frac{N\epsilon_{1}^{2}}{2B^{2}}\right)\right)\cdot\left(1-\exp\left(-\frac{N\epsilon_{2}^{2}}{2B^{2}}\right)\right)$ (29)
$=\left(1-\exp\left(-\frac{N\alpha^{2}\left({\mathbb{E}}[r(s_{q},a^{*})]-{\mathbb{E}}[r(s_{q},a)]\right)^{2}}{2B^{2}}\right)\right)\!\!\cdot\!\!\left(1-\exp\left(-\frac{N(1-\alpha)^{2}\left({\mathbb{E}}[r(s_{q},a^{*})]-{\mathbb{E}}[r(s_{q},a)]\right)^{2}}{2B^{2}}\right)\right).$ (30)
@@
$P\left(\frac{1}{N}\sum_{i=1}^{N}r(s_{q},a^{*}\mid\pi)\geq\frac{1}{N}\sum_{i=1}^{N}r(s_{q},a\mid\pi)\right)\!\geq\!\left(1-\exp\!\!\left(-\frac{N\left({\mathbb{E}}[r(s_{q},a^{*})]-{\mathbb{E}}[r(s_{q},a)]\right)^{2}}{8B^{2}}\right)\right)^{2}\!\!\!\!,\forall a\in A\setminus\{a^{*}\}.$ (31)
@@
$\bar{a}=\operatorname*{arg\,max}_{a\in A\setminus\{a^{*}\}}\frac{1}{N}\sum_{i=1}^{N}r(s_{q},a\mid\pi).$ (32)
@@
$P\left(\frac{1}{N}\sum_{i=1}^{N}r(s_{q},a^{*}\mid\pi)\geq\max_{a\in A\setminus\{a^{*}\}}\frac{1}{N}\sum_{i=1}^{N}r(s_{q},a\mid\pi)\right)$
@@
Hence, we require the trust horizon $N$ to satisfy
@@
B.2 Validation of Assumption 2 in the grid world MDP
For the sake of simplicity and without loss of generality, we consider a single-dimensional grid world MDP with the understanding that Assumption 2 holds for the two-dimensional grid world MDP as well, which is considered in our numerical experiments. The environmental details of the single-dimensional grid world MDP can be found in Figure 9. To proceed, we rely on the lemma below, and Assumption 2 is then validated by Proposition 1 that follows.
Figure 9: A single-dimensional grid world MDP comprising five states $\{s_{0},s_{1},s_{2},s_{3},s_{4}\}$, where $s_{0}$ represents the goal state (golden star). The environment offers two possible actions: $a_{0}$ that corresponds to moving left, and $a_{1}$ that corresponds to moving right. Crossing the boundaries is strictly prohibited. Any transitions that would result in boundary crossing will be confined to the current position. The reward structure is sparse, with a value of 1 received solely upon reaching the goal state $s_{0}$ and a value of 0 otherwise (see Figure 10 (left)). We consider an infinite time horizon with a discounter factor $\gamma$.
@@ B.2 Validation of Assumption 2 in the grid world MDP
With the understanding that the TD error ${\mathcal{E}}(s,a)$ of $Q_{\text{MDP}}^{\pi}(s,a)$ is zero, we iteratively learn and update $Q_{\text{MDP}}^{\pi}(s,a)$ by minimizing the TD error. To that end, we consider $\gamma=0.99$ and a convergence threshold $\epsilon_{Q}=10^{-6}$ (can be arbitrarily small). We initialize $Q_{\text{MDP}}^{\pi}(s,a)$ to be full of zeros as in Figure 10 (middle). Subsequently, we consistently update $Q_{\text{MDP}}^{\pi}(s,a)$ under the uniform policy $\pi$, until convergence as follows
@@
Empirically, we observe that the Q-table $Q_{\text{MDP}}^{\pi}(s,a)$ converges after the $1216^{th}$ iteration; demonstrated in Figure 11. As depicted in Figure 10 (right), the convergent $Q_{\text{MDP}}^{\pi}(s,a)$ implies that the optimal action for any state under the uniform policy $\pi$ is $a_{0}$ (see the golden stars), i.e.,
@@
Hence we obtain
@@
Figure 10: Left: reward table. Middle: initial Q-table. Right: convergent Q-table under uniform policy $\pi$.
@@
B.3 Proof of Theorem 2
Theorem 2 (MDP). Let Assumptions 1 and 2 hold. Denote by $a^{*}$ the optimal action given $s_{q}$. Consider a random policy $\pi$ as well as its Q-function $Q_{\text{MDP}}^{\pi}$. Consider the trust horizon $N>\log_{\gamma}\left(\kappa(1-\gamma)/(2B)\right)-1$. Define $\kappa=\min\limits_{s_{q}\in S}\left(Q_{\text{MDP}}^{\pi}(s_{q},a^{*})-\max\limits_{a\in A\setminus\{a^{*}\}}Q_{\text{MDP}}^{\pi}(s_{q},a)\right)$. With probability at least $1-\delta$, it holds that
$\hat{Q}_{\text{MDP}}^{\pi,N}(s_{q},a^{*})\geq\max_{a\in A\setminus\{a^{*}\}}\hat{Q}_{\text{MDP}}^{\pi,N}(s_{q},a),\,\forall s_{q}\in S,$ (75)
when the number of episodes $N_{\text{ep}}$ satisfies
$N_{\text{ep}}\geq\frac{2\left(1-\gamma^{N+1}\right)^{2}}{\left(\kappa\left(1-\gamma\right)/(2B)-\gamma^{N+1}\right)^{2}}\log\left(\frac{1+\sqrt{1-\delta}}{\delta}\right).$ (76)
@@ B.5 Proof of Corollary 1
B.6 Finite MDP setting from Osband et al. (2013)
@@ B.7 Proof of Corollary 2
Appendix C Implementation Details
@@
Algorithm 5 Collecting Query States and Action Labels under Random Policy (Dense-Reward MDP)
1:  Require: Random policy $\pi$, state space $S$, action space $A$, environment $\tau$, trust horizon $N$, empty return list $L_{r}$
2:  Sample a query state $s_{q}\sim S$
3:  for $a$ in $[A]$ do
4:     Initialize the state and action as $s_{0}=s_{q},\,a_{0}=a$
5:     Run an episode of $N$ steps in $\tau$ under the random policy $\pi$
6:     Add the discounted episodic return to $L_{r}$
7:  end for
8:  Obtain $a_{l}=A(\operatorname*{arg\,max}(L_{r}))$
9:  Return $(s_{q},a_{l})$
@@
Appendix D Experimental Details
D.1 Environmental Setup
Gaussian Bandits.
We investigate a five-armed bandit problem in which the state space $S$ consists solely of a singleton state $s_{q}$. With each arm (action) pulled, the agent receives a reward feedback. The goal is to identify the optimal arm that can maximize the cumulative reward. We consider the reward function for each arm following a Gaussian distribution with mean $\mu_{a}$ and variance $\sigma^{2}$, i.e., $R(\cdot|s_{q},a)={\mathcal{N}}(\mu_{a},\sigma^{2})$. Each arm possesses means $\mu_{a}$ drawn from a uniform distribution $U[0,1]$ and all arms share the same variance $\sigma=0.3$. We consider the pretraining and test data to have distinct Gaussian distributions with different means.
Bernoulli Bandits.
Darkroom.
Darkroom (Laskin et al., 2022; Zintgraf et al., 2019) is a two-dimensional navigation task with discrete state and action spaces. The room consists of $7\times 7$ grids ($|S|=49$), with an unknown goal randomly placed at any of these grids. The agent can select $5$ actions: go up, go down, go left, go right, or stay. The horizon length for Darkroom is $49$, meaning the agent must reach the goal within 49 moves. The challenge of this task arises from its sparse reward structure, i.e., the agent receives a reward of 1 solely upon reaching the goal, and 00 otherwise. Given $7\times 7=49$ available goals, we utilize $39$ of these goals ($\sim 80\%$) for pretraining and reserve the remaining 10 ($\sim 20\%$) (unseen during pretraining) for test.
Darkroom-Large.
Miniworld.
Miniworld is a three-dimensional pixel-based navigation task. The agent is situated in a room with four differently colored boxes, one of which is the target (unknown to the agent). The agent must navigate to the target box using $25\times 25\times 3$ image observations and by selecting from 4 available actions: turn left, turn right, move forward, or stay. Similar to Darkroom, the agent receives a reward of 1 only upon approaching the target box, and 0 otherwise. The high-dimensional pixel inputs clearly render Miniworld a much more challenging task than Darkroom and Darkroom-Large.
D.2 Hyperparameters

### v3

@@ 6 Conclusion
Appendix A Omitted Proofs
A.1 Proof of Theorem 1
Theorem 1 (MAB). Let Assumption 1 hold. The random policy is at least $(1-\delta)$-trustworthy as in Definition 1, when the trust horizon $N$ satisfies
@@ A.1 Proof of Theorem 1
$\frac{1}{N_{a^{*}}}\sum_{i=1}^{N_{a^{*}}}r(s_{q},a^{*})\geq{\mathbb{E}}[r(s_{q},a^{*})]-\epsilon_{1},$ (18)
$\frac{1}{N_{a}}\sum_{i=1}^{N_{a}}r(s_{q},a)\leq{\mathbb{E}}[r(s_{q},a)]+\epsilon_{2}.$ (19)
We note that the two inequalities above are the sufficient but not necessary conditions for $\frac{1}{N_{a^{*}}}\sum_{i=1}^{N_{a^{*}}}r(s_{q},a^{*})\geq\frac{1}{N_{a}}\sum_{i=1}^{N_{a}}r(s_{q},a)$ to hold.
@@
$P\left(\frac{1}{N_{a^{*}}}\sum_{i=1}^{N_{a^{*}}}r(s_{q},a^{*})\geq\frac{1}{N_{a}}\sum_{i=1}^{N_{a}}r(s_{q},a)\right)$
$\geq P\left(\frac{1}{N_{a^{*}}}\sum_{i=1}^{N_{a^{*}}}r(s_{q},a^{*})\geq{\mathbb{E}}[r(s_{q},a^{*})]-\epsilon_{1},\frac{1}{N_{a}}\sum_{i=1}^{N_{a}}r(s_{q},a)\leq{\mathbb{E}}[r(s_{q},a)]+\epsilon_{2}\right)$ (20)
$=P\left(\frac{1}{N_{a^{*}}}\sum_{i=1}^{N_{a^{*}}}r(s_{q},a^{*})\geq{\mathbb{E}}[r(s_{q},a^{*})]-\epsilon_{1}\right)\cdot P\left(\frac{1}{N_{a}}\sum_{i=1}^{N_{a}}r(s_{q},a)\leq{\mathbb{E}}[r(s_{q},a)]+\epsilon_{2}\right),$ (21)
@@
$P\left(\frac{1}{N_{a}}\sum_{i=1}^{N_{a}}r(s_{q},a)-{\mathbb{E}}[r(s_{q},a)]\leq\epsilon_{2}\right)\geq 1-\exp\left(-\frac{2N_{a}\epsilon_{2}^{2}}{(B-(-B))^{2}}\right),$ (22)
@@
$P\left(\frac{1}{N_{a}}\sum_{i=1}^{N_{a}}r(s_{q},a)-{\mathbb{E}}[r(s_{q},a)]\leq\epsilon_{2}\right)\geq 1-\exp\left(-\frac{N_{a}\epsilon_{2}^{2}}{2B^{2}}\right).$ (23)
@@
$P\left(\frac{1}{N_{a^{*}}}\sum_{i=1}^{N_{a^{*}}}r(s_{q},a^{*})-{\mathbb{E}}[r(s_{q},a^{*})]\geq-\epsilon_{1}\right)\geq 1-\exp\left(-\frac{2N_{a^{*}}\epsilon_{1}^{2}}{(B-(-B))^{2}}\right),$ (24)
@@
$P\left(\frac{1}{N_{a^{*}}}\sum_{i=1}^{N_{a^{*}}}r(s_{q},a^{*})-{\mathbb{E}}[r(s_{q},a^{*})]\geq-\epsilon_{1}\right)\geq 1-\exp\left(-\frac{N_{a^{*}}\epsilon_{1}^{2}}{2B^{2}}\right).$ (25)
Since $N\triangleq\min_{a\in A}N_{a}$, it then follows from the monotonicity of the exponential function that
$P\left(\frac{1}{N_{a^{*}}}\sum_{i=1}^{N_{a^{*}}}r(s_{q},a^{*})\geq\frac{1}{N_{a}}\sum_{i=1}^{N_{a}}r(s_{q},a)\right)$
$\geq\left(1-\exp\left(-\frac{N_{a^{*}}\epsilon_{1}^{2}}{2B^{2}}\right)\right)\cdot\left(1-\exp\left(-\frac{N_{a}\epsilon_{2}^{2}}{2B^{2}}\right)\right)$ (26)
$\geq\left(1-\exp\left(-\frac{N\epsilon_{1}^{2}}{2B^{2}}\right)\right)\cdot\left(1-\exp\left(-\frac{N\epsilon_{2}^{2}}{2B^{2}}\right)\right).$ (27)
@@
$P\left(\frac{1}{N_{a^{*}}}\sum_{i=1}^{N_{a^{*}}}r(s_{q},a^{*})\geq\frac{1}{N_{a}}\sum_{i=1}^{N_{a}}r(s_{q},a)\right)$
$\geq\left(1-\exp\left(-\frac{N\alpha^{2}\left({\mathbb{E}}[r(s_{q},a^{*})]-{\mathbb{E}}[r(s_{q},a)]\right)^{2}}{2B^{2}}\right)\right)\!\!\cdot\!\!\left(1-\exp\left(-\frac{N(1-\alpha)^{2}\left({\mathbb{E}}[r(s_{q},a^{*})]-{\mathbb{E}}[r(s_{q},a)]\right)^{2}}{2B^{2}}\right)\right).$ (28)
@@
$P\left(\frac{1}{N_{a^{*}}}\sum_{i=1}^{N_{a^{*}}}r(s_{q},a^{*})\geq\frac{1}{N_{a}}\sum_{i=1}^{N_{a}}r(s_{q},a)\right)\geq\left(1-\exp\!\left(-\frac{N\left({\mathbb{E}}[r(s_{q},a^{*})]-{\mathbb{E}}[r(s_{q},a)]\right)^{2}}{8B^{2}}\right)\right)^{2},\forall a\in A\setminus\{a^{*}\}.$ (29)
@@
$\bar{a}=\operatorname*{arg\,max}_{a\in A\setminus\{a^{*}\}}\frac{1}{N_{a}}\sum_{i=1}^{N_{a}}r(s_{q},a).$ (30)
@@
$P\left(\frac{1}{N_{a^{*}}}\sum_{i=1}^{N_{a^{*}}}r(s_{q},a^{*})\geq\max_{a\in A\setminus\{a^{*}\}}\frac{1}{N_{a}}\sum_{i=1}^{N_{a}}r(s_{q},a)\right)$
@@
Hence, we require the trust horizon $N\triangleq\min_{a\in A}N_{a}$ to satisfy
@@
A.2 Validity of Assumption 2 in the grid world MDP with a single sparse reward
For the sake of simplicity and without loss of generality, we consider a single-dimensional grid world MDP with the understanding that Assumption 2 holds for the two-dimensional grid world MDP as well, which is considered in our numerical experiments. The environmental details of the single-dimensional grid world MDP can be found in Figure 2, where the reward is received only upon reaching the unique goal. To proceed, we rely on the lemma below, and Assumption 2 is then validated by Proposition 1 that follows.
@@ A.2 Validity of Assumption 2 in the grid world MDP with a single sparse reward
Consider the MDP of a single-dimensional grid world with $S=\{s_{0},s_{1},s_{2},s_{3},s_{4}\}$ and $A=\{a_{0},a_{1}\}$, as depicted in Figure 2. Consider a uniform random policy $\pi$. It holds that
@@
With the understanding that the TD error ${\mathcal{E}}(s,a)$ of $Q_{\text{MDP}}^{\pi}(s,a)$ is zero, we iteratively learn and update $Q_{\text{MDP}}^{\pi}(s,a)$ by minimizing the TD error. To that end, we consider $\gamma=0.99$ and a convergence threshold $\epsilon_{Q}=10^{-6}$ (can be arbitrarily small). We initialize $Q_{\text{MDP}}^{\pi}(s,a)$ to be full of zeros as in Figure 9 (middle-left). Subsequently, we consistently update $Q_{\text{MDP}}^{\pi}(s,a)$ under the uniform policy $\pi$, until convergence as follows
@@
Empirically, we observe that the Q-table $Q_{\text{MDP}}^{\pi}(s,a)$ converges after the $1216^{th}$ iteration; demonstrated in Figure 10. As depicted in Figure 9 (middle-right), the convergent $Q_{\text{MDP}}^{\pi}(s,a)$ implies that the optimal action for any state under the uniform policy $\pi$ is $a_{0}$ (see the golden stars), i.e.,
@@
Likewise, we follow the same process above using the Bellman optimality equation
$Q_{\text{MDP}}^{*}(s,a)=r(s,a)+\gamma\max_{a^{\prime}\in A}\,Q_{\text{MDP}}^{*}(s^{\prime},a^{\prime}),\,\forall(s,a)\in S\times A.$ (72)
The convergent $Q_{\text{MDP}}^{*}(s,a)$ in Figure 9 (right) demonstrates that the optimal action for any state under the optimal policy is $a_{0}$ as well, i.e.,
$\operatorname*{arg\,max}_{a\in A}\,Q_{\text{MDP}}^{*}(s,a)=a_{0},\,\forall s\in S.$ (73)
@@
Figure 9: Left: reward table. Middle-Left: initial Q-table. Middle-Right: convergent Q-table under uniform policy $\pi$. Right: convergent optimal Q-table.
@@
A.3 Discussion of Assumption 2 in the grid world MDP with two sparse rewards
Figure 11: A single-dimensional grid world MDP comprising five states $\{s_{0},s_{1},s_{2},s_{3},s_{4}\}$. The environment offers two possible actions: $a_{0}$ that corresponds to moving left, and $a_{1}$ that corresponds to moving right. Crossing the boundaries is strictly prohibited. Any transitions that would result in boundary crossing will be confined to the current position. The reward structure is sparse, with a value of $R_{l}$ received upon reaching $s_{0}$, a value of $R_{s}$ received upon reaching $s_{4}$ and a value of 0 otherwise. We consider an infinite time horizon with a discounter factor $\gamma$.
As depicted in Figure 11, we consider the same problem setting as in Appendix A.2 except that we assign the rewards $R_{l}$ and $R_{s}$ to the states $s_{0}$ and $s_{4}$, respectively. In the following three cases, we demonstrate that Assumption 2 holds for some (but not all) navigation problems that have two sparse rewards. The validity of Assumption 2 depends on the problem-specific factors such as the rewards and the discount factor $\gamma$.
  • (a)
    Consider $R_{l}=1,R_{s}=0.1,\gamma=0.99$. The action that maximizes the Q table for both random and optimal policies is the same, i.e., $a_{0}=\arg\max_{a}\,Q_{MDP}^{\pi}(s,a)=\arg\max_{a}\,Q_{MDP}^{*}(s,a)$ for all states as shown in Figure 12.
    Figure 12: When $R_{l}=1,R_{s}=0.1,\gamma=0.99$. Left: convergent Q-table under uniform policy $\pi$. Right: convergent optimal Q-table.
  • (b)
    Consider $R_{l}=1,R_{s}=0.9,\gamma=0.8$. In this case, the random policy selects $a_{0}$ at the states $s_{0},s_{1},s_{2}$ and $a_{1}$ at the states $s_{3},s_{4}$ (see $Q_{MDP}^{\pi}$ in Figure 13). However, notice that the optimal policy selects exactly the same optimal actions as that of the random policy, i.e., $a_{0}$ for states $s_{0},s_{1},s_{2}$ and $a_{1}$ for states $s_{3},s_{4}$ (see $Q_{MDP}^{*}$ in Figure 13), thus Assumption 2 still holding.
    Figure 13: When $R_{l}=1,R_{s}=0.9,\gamma=0.8$. Left: convergent Q-table under uniform policy $\pi$. Right: convergent optimal Q-table.
  • (c)
    Consider $R_{l}=1,R_{s}=0.9,\gamma=0.97$. The convergent Q-tables in Figure 14 indicate that under the random policy, the optimal action for states $s_{3}$ and $s_{4}$ is $a_{1}$, whereas under the optimal policy, it is $a_{0}$. This discrepancy indicates that Assumption 2 does not hold.
    Figure 14: When $R_{l}=1,R_{s}=0.9,\gamma=0.97$. Left: convergent Q-table under uniform policy $\pi$. Right: convergent optimal Q-table.
A.4 Proof of Theorem 2
Theorem 2 (MDP). Let Assumptions 1 and 2 hold. Define $\kappa=\min\limits_{s_{q}\in S}\big{(}Q_{\text{MDP}}^{\pi}(s_{q},a^{*})-\max\limits_{a\in A\setminus\{a^{*}\}}Q_{\text{MDP}}^{\pi}(s_{q},a)\big{)}$. Consider the trust horizon $N>\log_{\gamma}\left(\kappa(1-\gamma)/(2B)\right)-1$. The random policy is at least $(1-\delta)$-trustworthy as in Definition 2, when the number of episodes $N_{\text{ep}}$ satisfies
@@ A.6 Proof of Corollary 1
A.7 Finite MDP setting from Osband et al. (2013)
@@ A.8 Proof of Corollary 2
Appendix B Implementation Details
@@
[…]
4:     Sample a query state $s_{q}\sim S$
5:     Empty a step list $L_{s}$
6:     for $a$ in $[A]$ do
7:        Initialize the state and action as $s_{0}=s_{q},\,a_{0}=a$
9:        Add consumed steps to $L_{s}$ (add “$N+1$” if no reward is received)
10:     end for
[…]
12:  end while
13:  Obtain $a_{l}=A(\operatorname*{arg\,min}(L_{s}))$
14:  Return $(s_{q},a_{l})$
@@
Appendix C Experimental Details
C.1 Hyperparameters